OpenStudy (anonymous):

the solid lies between planes perpendicular to the x axis at x= - 4 and x = 4. The cross sections perpendicular to the x axis between these planes are squares whose diagonals run from the semicircle y = -sqrt(16 - x^2) to the semicircle y = sqrt(16 - x^2). Find the volume of the described solid.

OpenStudy (anonymous):

AP Calculus AB question

OpenStudy (anonymous):

I got 512/3, can anyone verify?

OpenStudy (anonymous):



OpenStudy (anonymous):

did I draw it right?

OpenStudy (anonymous):



OpenStudy (anonymous):

okay yes second drawing is what I envisioned, but there isn't a "correct" drawing provided if that's what you're asking

OpenStudy (anonymous):

I'll answer in 15 minutes, I have to drive home now :DD

OpenStudy (anonymous):

thanks both of you! ganeshie8 I ended up getting 2(16-x^2) as the area formula and integrating it from -4 to 4 as well. The problem is is that this is a multiple choice question and 512/3 is not an answer choice

ganeshie8 (ganeshie8):

^ same answer you did \(\int (side)^2 dx\) righ ?t

ganeshie8 (ganeshie8):

okay lets go thru it again

OpenStudy (anonymous):

Yes, ∫(side)2dx, given that the side was the diagonal squared times 1/2 and the diagonal was the distance between the functions\[\sqrt{16-x ^{2}}\] and \[-\sqrt{16-x ^{2}}\]

ganeshie8 (ganeshie8):

we're on same page

OpenStudy (anonymous):

i think this chalks up to a teacher error, especially since this wasn't the only problem I found with the multiple choice answers. I can't think of a single other way to do the problem